A mass of a gas occupies 20 cm^3 at 5 degree celcius and 760 mm Hg pressure, what is its approximate volume at 30 degree celcius and 800 mm Hg pressure?
1
Expert's answer
2020-07-03T10:00:55-0400

According to the general gas equation, the volume "V", temperature "T" and pressure "P" of the ideal gas are connected in the following way:


"\dfrac{PV}{T} = const"

Thus, denoting with the index 1 parameters for the initial state and with the index 2 parameters for the final state, obtain:

"\dfrac{P_1V_1}{T_1} = \dfrac{P_2V_2}{T_2}"

Expressing the final volume from this equation, get:


"V_2 = \dfrac{P_1V_1T_2}{P_2T_1}"

Substituting the numerical value (and converting the temperature to kelvin), get:


"V_2 = \dfrac{760\cdot 20\cdot (273+5)}{800\cdot (273+30)} \approx 17.4 cm^3"

Answer. 17.4 cm^3.